SCIENTIFIC NOTATION (1.23 x 10^n).


Three point charges are arranged on a line. Charge q3 = +5.00 nC and is at the origin. Charge q2 = -3.00 nC and is at x = +4.00 cm. Charge q1 is at x = +2.00 cm. What is q1 (magnitude and sign) if the net force on q3 is zero? 






1
Expert's answer
2021-05-31T08:39:03-0400

"F_{23}=-F_{13}\to k\frac{q_2q_3}{0.04^2}=-k\frac{q_1q_3}{0.02^2}"


"\frac{q_2}{0.04^2}=-\frac{q_1}{0.02^2}\to q_1=-\frac{-3\cdot10^{-9}}{0.04^2}\cdot0.02^2=+0.75\cdot10^{-9}\ (C)="



"=+0.75\ (nC)" . Answer